a uniform rope of length L and mass m is pulled by a contact force F. find the tension in the rope at a distance x from the end where it is applied.

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Given : Lenght of Rope = L
Mass of Rope = m
Tension Of Rope= T at x distance from end of rope
Acceleration = a
The the linear density of Rope = M/L
Mass = Linear Density x length
Mass of x Meters rope = Mx/L
Mass of (L-x) Meter Rope = {m(L-x)}/L 
System I : T₂ = [{m(L-x)}/L](a)
 
Regards
Arun (askIITians forum expert)